# Thresholds for the distributed surface code in the presence of memory decoherence

### v1

@@ I Introduction
A distributed quantum computer ciracDistributedQuantumComputation1999 ; groverQuantumTelecomputation1997 realizes a large-scale processing system by using entanglement to link smaller quantum processing units. For example, the sub-units may be elements of a photonic chip, or form the nodes of a quantum network on a larger scale. vanmeterPathScalableDistributed2016 Fault tolerance is naturally achieved by establishing the connectivity according to the architecture of a topological error-correction code. The distributed approach provides advantages in terms of scalability but is limited by the availability of high-quality entanglement. This is because entangled states are required for both inter-node operations and for detecting local errors with the error-correction code. fujiiDistributedArchitectureScalable2012 ; liHighThresholdDistributed2012a ; monroeLargeScaleModular2014
@@
In this paper, we focus on systems that are capable of generating remote two-qubit entanglement between pairs of connected nodes. There exist several physical systems suitable for generating this type of entanglement with optical interfaces. awschalomQuantumTechnologiesOptically2018 Examples of this are ion traps, neutral atoms, and color centers in the diamond lattice, such as nitrogen-vacancy (NV), bernienHeraldedEntanglementSolidstate2013 ; Taminiau2014 ; riedelDeterministicEnhancementCoherent2017 ; humphreysDeterministicDeliveryRemote2018 ; Bradley2019 ; Abobeih2019 ; rufResonantExcitationPurcell2021 ; pompiliRealizationMultinodeQuantum2021 ; abobeihFaulttolerantOperationLogical2022 ; bradleyRobustQuantumnetworkMemory2022 ; orphal-kobinOpticallyCoherentNitrogenVacancy2023 silicon-vacancy (SiV), Sipahigil2016 ; sukachevSiliconVacancySpinQubit2017 ; nguyenIntegratedNanophotonicQuantum2019 ; nguyenQuantumNetworkNodes2019 ; knautEntanglementNanophotonicQuantum2023 or tin-vacancy (SnV) iwasakiTinVacancyQuantumEmitters2017 ; rugarQuantumPhotonicInterface2021 ; debrouxQuantumControlTinVacancy2021 centers. As a concrete example, we investigate the distributed surface code with hardware-specific noise based on color centers, also known as defect centers. We emphasize that the obtained insights are more general and that our simulation tools allow for implementing error models based on general hardware implementations.
Diamond color centers host long-lived electron spins that exhibit coherent optical transitions, enabling their use as a communication qubit. This qubit is used to create entanglement with other nodes and further can address up to tens of proximal nuclear spins and or other electron spins occurring in the host material. humphreysDeterministicDeliveryRemote2018 ; pompiliRealizationMultinodeQuantum2021 ; knautEntanglementNanophotonicQuantum2023 These nuclear spins can be used as local processing qubits—i.e., as a local memory to store and manipulate quantum states. Bradley2019 Hereafter, in the context of diamond color centers, the term “memory qubits” specifically refers to such spins.
@@
The toric code has a depolarizing phenomenological error probability threshold of approximately 10% to 11%.brownedanLecturesTopologicalCodes This error model assumes that all qubits of the code are part of the same quantum system, stabilizer measurements can be carried out perfectly, and the qubits experience depolarizing noise in between the stabilizer measurement rounds. A more precise analysis with a circuit-level error model yields error probability thresholds between 0.9% and 0.95%.Nickerson2013 In this model, the stabilizer measurement circuit is simulated with noisy gates and measurements and it is implicitly assumed that the connectivity of the system allows direct two-qubit gates between adjacent qubits of the code topology. Therefore, this error model corresponds to a monolithic architecture.
@@ II.2 Dynamic program to generate GHZ protocols
Fusion is executed to create GHZ states out of Bell pairs and to create a larger GHZ state. A state of the form |$\ket{\text{GHZ}_{n_{1}}}$ can be fused with a state |$\ket{\text{GHZ}_{n_{2}}}$ by applying a CNOT gate between one qubit of both states and measuring out the target qubit in the $Z$ basis. Obtaining a $+1$ measurement outcome results in the state |$\ket{\text{GHZ}_{n_{1}+n_{2}-1}}$. A $-1$ measurement outcome leads to the same state after local Pauli-$X$ corrections.
@@
In Alg. II.2, we present a schematic, pseudo-code version of the dynamic program we used to generate and evaluate GHZ protocols. This algorithm is an expanded version of the dynamic program in Ref. deboneProtocolsCreatingDistilling2020, . In this algorithm, each protocol is created with either a fusion or a distillation operation that combines two smaller GHZ protocols encountered earlier in the search. The protocols created in this fashion can be depicted with a directed binary tree graph. An example graph is given on the left side of Fig. 3. For the distillation steps in the binary tree diagrams, we consume the state on the right to purify the state on the left.
@@
We considered two heuristics to limit the influence of decoherence and of SWAP gates. To limit decoherence, we prioritize creating larger branches of the tree. Here, a branch is defined as an element of the binary tree (i.e., an operation in the GHZ protocol) including all elements that (in)directly point towards it. The size of the branch is the number of elements it contains. Because, generally speaking, creating small branches is faster than creating large branches, this heuristic aims to minimize waiting times for completed intermediate branches of the GHZ protocol.
@@
Finally, while this dynamic program explores a large class of protocols, not all of the seven protocols that we introduced in the previous section can be generated. This is because, to suppress calculation time, the program limits distillation steps to operations that use an ancillary entangled state to non-locally measure a stabilizer operator of the main state, in a sequential manner. The protocols Refined, Expedient and Stringent, however, make use of the so-called double selection distillation block fujiiEntanglementPurificationDouble2009 that does not directly fit into this stabilizer distillation framework.
Figure 3: (a) Binary tree with $k=7$ found with the dynamic program of Alg. II.2: the Septimum protocol. In this directed graph, the top vertex represents the final state. Each vertex describes how its corresponding state is created from a fusion or distillation operation involving its two children. At the origin of each branch, we find the elementary links: the Bell pairs. (b) We split the binary tree into multiple time steps that describe the order in which the protocol is carried out. The subtree involving the links between nodes $C$ and $D$ is identified as the part that we want to carry out first; the subtree involving links between $A$ and $B$ is also added to time step 1, because it can be carried out in parallel. (c) The timed binary tree is converted to an explicit set of operations: a protocol recipe. Here, we also add necessary SWAP gates, conditional corrections for fusion operations, and evaluations of distillation operations. For distillation operations, we also add instructions in case of failure (not printed here). During the execution of this protocol, the system waits until all branches of a time step are completed before continuing to the next time step. This protocol recipe uses a maximum of $q=3$ qubits per network node to generate the GHZ state.
@@ III.1 The toric surface code
The measurement outcomes of stabilizers, known as the error syndrome, are fed to a decoder to estimate the underlying error pattern. Here, we consider an implementation by Hu huQuasilinearTimeDecoding2020 ; huOOPSurfaceCode of the Union-Find delfosseAlmostlinearTimeDecoding2021 error decoder.
@@ III.2 Distributed toric code simulation
Due to entanglement distillation steps in the GHZ creation protocol, GHZ generation is probabilistic. In order to fix the duration of the rounds we impose a “GHZ cycle time” $t_{\mathrm{GHZ}}$: if GHZ generation is not successful within this time, it aborts. In that case, the corresponding stabilizer operator can not be measured. This information could be given to the decoder in the form of an erasure symbol. However, to leverage existing decoders, we opt to duplicate the last measured value of the stabilizer. This choice is suboptimal and better thresholds could be expected for decoders that can handle erasures and noisy measurements. The GHZ cycle time is a free parameter that we explore in Sec. V. In App. B.5, we describe a heuristic-driven approach for selecting a suitable GHZ cycle time.
@@ IV Diamond color center model
We assume that all operations take a finite amount of time. The time durations can be found in Table 1. Furthermore, we assume single-qubit gates to be noiseless, while noise in two-qubit gates is modeled with a depolarizing channel with probability $p_{\text{g}}$ (see App. B.4). We model imperfect measurements by flipping the measurement outcome with probability $p_{\text{m}}$.
@@ V Results
Protocol
Single-click cabrilloCreationEntangledStates1999
Double-click barrettEfficientHighfidelityQuantum2005a
$F_{\text{prep}}$
$0.99$hensenLoopholefreeBellInequality2015a ; rozpedekNeartermQuantumrepeaterExperiments2019
$0.999$
$p_{\text{EE}}$
$0.04$humphreysDeterministicDeliveryRemote2018
$p_{\text{EE}}(f_{\phi})$
$0.01$
$\mu$
$0.9$humphreysDeterministicDeliveryRemote2018 ; pompiliRealizationMultinodeQuantum2021
$0.95$
$\lambda$
$0.984$humphreysDeterministicDeliveryRemote2018 ; dahlbergLinkLayerProtocol2019
$1$
$\eta_{\mathrm{ph}}$
$0.0046$coopmansNetSquidNETworkSimulator2021
$0.4472$
$\eta_{\mathrm{ph}}(f_{\eta})$
@@
$p_{\text{link}}$
$0.0001$
$0.1$
$p_{\text{link}}(f_{\eta})$
$F_{\text{link}}$
$0.8966$
$F_{\text{link}}(f_{\phi})$
$0.9526$
@@
$t_{\text{link}}$
$6\cdot 10^{-6}$ s
$t_{\text{meas}}$
$4\cdot 10^{-6}$ s
$t_{X,Y}^{\text{e}}$
$0.14\cdot 10^{-6}$ s
$t_{X,Y}^{\text{n}}$
$1.0\cdot 10^{-3}$ s
$t_{Z,H}^{\text{e}}$
$0.1\cdot 10^{-6}$ s
$t_{Z,H}^{\text{n}}$
$0.5\cdot 10^{-3}$ s
$t_{\text{C}Z,\text{C}X,\text{C}iY}$
$0.5\cdot 10^{-3}$ s
$t_{\text{SWAP}}$
$1.5\cdot 10^{-3}$ s
@@
$T1_{\text{idle}}^{\text{n}}$
$300$ s
$T1_{\text{link}}^{\text{n}}$
$0.03$ s pompiliRealizationMultinodeQuantum2021
$0.3$ s
$0.03f_{\text{dec}}$ s
$0.3$ s
$T1_{\text{idle}}^{\text{e}}$
$300$ s
$T2_{\text{idle}}^{\text{n}}$
$10$ s
$T2_{\text{link}}^{\text{n}}$
$0.0075$ s pompiliRealizationMultinodeQuantum2021
$0.075$ s
$0.0075f_{\text{dec}}$ s
$0.075$ s
$T2_{\text{idle}}^{\text{e}}$
$1.0$ s
$t_{\text{pulse}}$
$1.0\cdot 10^{-3}$ s
$n_{\text{DD}}$
500
18
Eq. (8)
@@
$\eta_{\mathrm{link}}^{*}$
$2\cdot 10^{-1}$
$2\cdot 10^{3}$
$200f_{\mathrm{dec}}$
$\eta_{\mathrm{link}}^{*}(f_{\eta})$
@@
$p_{\text{g}}$
$0.01$
Threshold
$p_{\text{m}}$
$0.01$
@@ V.1 Current state-of-the-art parameter set
Let us first consider a parameter set inspired by state-of-the-art NV center hardware (see the first column of Table 1). The operation times in this set are based on typical time scales in nitrogen-vacancy centers with a natural ${}^{13}$ concentration Bradley2019 ; pompiliRealizationMultinodeQuantum2021 ; bradleyRobustQuantumnetworkMemory2022 —see App. A for more details. The Bell pair parameter values are a collection of the best parameters in current NV center literature—see the first column of Table 1 for relevant citations. In the following, we explicitly refer to this parameter set as the “state-of-the-art” parameter set. As discussed in more detail in App. A, for this set, the single-click entanglement protocol outperforms the double-click protocol.
@@ V.2.1 Sensitivity to Bell pair fidelity
Firstly, we investigate the influence of the Bell pair fidelity by using a near-future setting parameter set—see the second column in Table 1. Compared to the state-of-the-art parameter set of Sec. V.1, in this set coherence times during entanglement creation and the photon detection probability are one and two orders of magnitude higher, respectively. The double-click entanglement protocol now gives rise to best combination of entanglement success probability and Bell pair fidelity, as explained in more detail in App. A. This makes that these near-future parameters allow for an increase in the link efficiency by four orders of magnitude compared to the state-of-the-art parameter set of Sec. V.1—see Eqs. (1), (3), and (5).
@@
In our Bell pair model, several parameters contribute to the infidelity of the Bell pair states in a similar way—i.e., through the parameter $\phi$ of Eq. (4) that captures all dephasing noise of the model. To investigate the sensitivity of the performance with respect to the Bell pair fidelity, we vary the influence of dephasing by scaling the probability of double excitation probability and off-resonant excitation errors. These are considered one of the leading error sources in present experiments. hermansQubitTeleportationNonneighbouring2022 We show the results in Fig. 6. In this figure, the bottom of the horizontal axis indicates the Bell pair fidelity; the top indicates the corresponding excitation error probability.

### v2

@@ I Introduction
A distributed quantum computer groverQuantumTelecomputation1997 ; ciracDistributedQuantumComputation1999 realizes a large-scale processing system by using entanglement to link smaller quantum processing units. For example, the sub-units may be elements of a photonic chip, or form the nodes of a quantum network on a larger scale. vanmeterPathScalableDistributed2016 Fault tolerance is naturally achieved by establishing the connectivity according to the architecture of a topological error-correction code. The distributed approach provides advantages in terms of scalability but is limited by the availability of high-quality entanglement. This is because entangled states are required for both inter-node operations and for detecting local errors with the error-correction code. jiangDistributedQuantumComputation2007 ; vanmeterDistributedQuantumComputation2010 ; fujiiDistributedArchitectureScalable2012 ; liHighThresholdDistributed2012a ; monroeLargeScaleModular2014 ; rametteFaulttolerantConnectionErrorcorrected2023
@@
In this paper, we focus on systems that are capable of generating remote two-qubit entanglement between pairs of connected nodes. There exist several physical systems suitable for generating this type of entanglement with optical interfaces. awschalomQuantumTechnologiesOptically2018 Examples of this are ion traps, neutral atoms, and color centers in the diamond lattice, such as nitrogen-vacancy (NV), bernienHeraldedEntanglementSolidstate2013 ; Taminiau2014 ; riedelDeterministicEnhancementCoherent2017 ; Cramer2016 ; humphreysDeterministicDeliveryRemote2018 ; Bradley2019 ; Abobeih2019 ; rufResonantExcitationPurcell2021 ; pompiliRealizationMultinodeQuantum2021 ; abobeihFaulttolerantOperationLogical2022 ; bradleyRobustQuantumnetworkMemory2022 ; orphal-kobinOpticallyCoherentNitrogenVacancy2023 silicon-vacancy (SiV), Sipahigil2016 ; sukachevSiliconVacancySpinQubit2017 ; nguyenIntegratedNanophotonicQuantum2019 ; nguyenQuantumNetworkNodes2019 ; knautEntanglementNanophotonicQuantum2023 or tin-vacancy (SnV) iwasakiTinVacancyQuantumEmitters2017 ; rugarQuantumPhotonicInterface2021 ; debrouxQuantumControlTinVacancy2021 centers. As a concrete example, we investigate the distributed surface code with hardware-specific noise based on color centers, also known as defect centers. We emphasize that the obtained insights are more general and that our simulation tools allow for implementing error models based on general hardware implementations.
Diamond color centers host long-lived electron spins that exhibit coherent optical transitions, enabling their use as a communication qubit. This qubit is used to create entanglement with other nodes and can address up to tens of proximal nuclear spins and or other electron spins occurring in the host material. humphreysDeterministicDeliveryRemote2018 ; pompiliRealizationMultinodeQuantum2021 ; knautEntanglementNanophotonicQuantum2023 These nuclear spins can be used as local processing qubits—i.e., as a local memory to store and manipulate quantum states. Bradley2019 Hereafter, in the context of diamond color centers, the term “memory qubits” specifically refers to such spins.
@@
The toric code has a depolarizing phenomenological error probability threshold of approximately 10% to 11%.brownedanLecturesTopologicalCodes2014 This error model assumes that all qubits of the code are part of the same quantum system, stabilizer measurements can be carried out perfectly, and the qubits experience depolarizing noise in between the stabilizer measurement rounds. A more precise analysis with a circuit-level error model yields error probability thresholds between 0.9% and 0.95%.Nickerson2013 In this model, the stabilizer measurement circuit is simulated with noisy gates and measurements and it is implicitly assumed that the connectivity of the system allows direct two-qubit gates between adjacent qubits of the code topology. Therefore, this error model corresponds to a monolithic architecture.
@@ II.2 Dynamic program to generate GHZ protocols
Fusion is executed to create GHZ states out of Bell pairs and to create a larger GHZ state. A state of the form |$\ket{\text{GHZ}_{n_{1}}}$ can be fused with a state |$\ket{\text{GHZ}_{n_{2}}}$ by applying a CNOT gate between one qubit of both states and measuring out the target qubit in the Pauli-$Z$ basis. Obtaining a $+1$ measurement outcome results in the state |$\ket{\text{GHZ}_{n_{1}+n_{2}-1}}$. A $-1$ measurement outcome leads to the same state after local Pauli-$X$ corrections.
@@
In Alg. II.2, we present a schematic, pseudo-code version of the dynamic program we used to generate and evaluate GHZ protocols. This algorithm is an expanded version of the dynamic program in Ref. deboneProtocolsCreatingDistilling2020, . In this algorithm, each protocol is created with either a fusion or a distillation operation that combines two smaller GHZ protocols encountered earlier in the search. The protocols created in this fashion can be depicted with a directed binary tree graph. An example graph is given on the left side of Fig. 3. For the distillation steps in the binary tree diagrams, we consume the state on the right to distill the state on the left.
@@
Whereas we did not optimize over the conversion from binary tree to protocol recipe, we considered two heuristics to limit the influence of decoherence and of SWAP gates. To limit decoherence, we prioritize creating larger branches of the tree. Here, a branch is defined as an element of the binary tree (i.e., an operation in the GHZ protocol) including all elements that (in)directly point towards it. The size of the branch is the number of elements it contains. Because, generally speaking, creating small branches is faster than creating large branches, this heuristic aims to minimize waiting times for completed intermediate branches of the GHZ protocol.
@@
In constructing the protocol recipe, we first use these heuristics to determine the order in which the elementary Bell pairs are generated—i.e., the leaves of the binary tree. By following this order, we then check for each Bell pair if other Bell pairs in non-overlapping network nodes can be generated simultaneously. Here, we prioritize based on proximity in the binary tree. We include instructions for distillation, fusion, and SWAP operations at the earliest possible point of execution. This approach gives rise to a unique conversion from binary tree to protocol recipe. More detailed descriptions of the protocol recipe construction and execution procedure can be found in Ref. deboneGHZDistillationProtocols2023, and the supplementary documents of the repository of Ref. deboneDataSoftwareUnderlying2024, .
While our dynamic program explores a large class of protocols, not all of the seven protocols that we introduced in the previous section can be generated. This is because, to suppress calculation time, the program limits distillation steps to operations that use an ancillary entangled state to non-locally measure a stabilizer operator of the main state, in a sequential manner. The protocols Refined, Expedient and Stringent, however, make use of the so-called double selection distillation block fujiiEntanglementPurificationDouble2009 that does not directly fit into this stabilizer distillation framework. krastanovOptimizedEntanglementPurification2019 ; jansenEnumeratingAllBilocal2022 ; goodenoughNeartermDistillationProtocols2024
Figure 3: (a) Binary tree with $k=7$ found with the dynamic program of Alg. II.2: the Septimum protocol. In this directed graph, the top vertex represents the final state. Each vertex describes how its corresponding state is created from a fusion or distillation operation involving its two children. At the origin of each branch, we find the elementary links: the Bell pairs. (b) We split the binary tree into multiple time steps that describe the order in which the protocol is carried out. The subtree involving the links between nodes $C$ and $D$ is identified as the part that we want to carry out first, since it is the left part of the largest branch of the binary tree. The subtree involving links between $A$ and $B$ is also added to time step 1, because it can be carried out in parallel. (See Sec. II.2 for more information.) (c) The timed binary tree is converted to an explicit set of operations: a protocol recipe. Here, we also add necessary SWAP gates, conditional corrections for fusion operations, and evaluations of distillation operations. For distillation operations, we also add instructions in case of failure (not printed here). During the execution of this protocol, the system waits until all branches of a time step are completed before continuing to the next time step. This protocol recipe uses a maximum of $q=3$ qubits per network node to generate the GHZ state.
@@ III.1 The toric surface code
The outcomes of the stabilizer measurements, known as the error syndrome, are fed to a decoder to estimate the underlying error pattern. Here, we consider an implementation by Hu huQuasilinearTimeDecoding2020 ; huOOPSurfaceCode of the Union-Find delfosseAlmostlinearTimeDecoding2021 error decoder.
@@ III.2 Distributed toric code simulation
Due to entanglement distillation steps in the GHZ creation protocol, GHZ generation is probabilistic. To fix the duration of the rounds we impose a “GHZ cycle time” $t_{\mathrm{GHZ}}$: if GHZ generation is not successful within this time, it aborts. In that case, the corresponding stabilizer operator can not be measured. This information could be given to the decoder in the form of an erasure symbol. However, to leverage existing decoders, we opt to duplicate the last measured value of the stabilizer. This choice is suboptimal and better thresholds could be expected for decoders that can handle erasures and noisy measurements. The GHZ cycle time is a free parameter that we explore in Sec. V. In App. B.5, we describe a heuristic-driven approach for selecting a suitable GHZ cycle time.
@@ IV Diamond color center model
We assume that all operations take a finite amount of time. The time durations can be found in Table 1. We neglect the influence of classical communication times, as we consider distances between network nodes to be relatively small, but include synchronization of network nodes when classical communication is required. Furthermore, we assume single-qubit gates to be noiseless, while noise in two-qubit gates is modeled with a depolarizing channel with probability $p_{\text{g}}$ (see App. B.4). We model imperfect measurements by flipping the measurement outcome with probability $p_{\text{m}}$.
@@ V Results
Protocol Single-click cabrilloCreationEntangledStates1999 Double-click barrettEfficientHighfidelityQuantum2005a
$F_{\text{prep}}$ $0.99$hensenLoopholefreeBellInequality2015a ; rozpedekNeartermQuantumrepeaterExperiments2019 $0.999$
$p_{\text{EE}}$ $0.04$humphreysDeterministicDeliveryRemote2018 $p_{\text{EE}}(f_{\phi})$ $0.01$
$\mu$ $0.9$humphreysDeterministicDeliveryRemote2018 ; pompiliRealizationMultinodeQuantum2021 $0.95$
$\lambda$ $0.984$humphreysDeterministicDeliveryRemote2018 ; dahlbergLinkLayerProtocol2019 $1$
$\eta_{\mathrm{ph}}$ $0.0046$coopmansNetSquidNETworkSimulator2021 $0.4472$ $\eta_{\mathrm{ph}}(f_{\eta})$
@@
$p_{\text{link}}$ $0.0001$ $0.1$ $p_{\text{link}}(f_{\eta})$
$F_{\text{link}}$ $0.8966$ $F_{\text{link}}(f_{\phi})$ $0.9526$
@@
$t_{\text{link}}$ $6\cdot 10^{-6}$ s
$t_{\text{meas}}$ $4\cdot 10^{-6}$ s
$t_{X,Y}^{\text{e}}$ $0.14\cdot 10^{-6}$ s
$t_{X,Y}^{\text{n}}$ $1.0\cdot 10^{-3}$ s
$t_{Z,H}^{\text{e}}$ $0.1\cdot 10^{-6}$ s
$t_{Z,H}^{\text{n}}$ $0.5\cdot 10^{-3}$ s
$t_{\text{C}Z,\text{C}X,\text{C}iY}$ $0.5\cdot 10^{-3}$ s
$t_{\text{SWAP}}$ $1.5\cdot 10^{-3}$ s
@@
$T1_{\text{idle}}^{\text{n}}$ $300$ s
$T1_{\text{link}}^{\text{n}}$ $0.03$ s pompiliRealizationMultinodeQuantum2021 $0.3$ s $0.03f_{\text{dec}}$ s $0.3$ s
$T1_{\text{idle}}^{\text{e}}$ $300$ s
$T2_{\text{idle}}^{\text{n}}$ $10$ s
$T2_{\text{link}}^{\text{n}}$ $0.0075$ s pompiliRealizationMultinodeQuantum2021 $0.075$ s $0.0075f_{\text{dec}}$ s $0.075$ s
$T2_{\text{idle}}^{\text{e}}$ $1.0$ s
$t_{\text{pulse}}$ $1.0\cdot 10^{-3}$ s
$n_{\text{DD}}$ 500 18 Eq. (8)
@@
$\eta_{\mathrm{link}}^{*}$ $2\cdot 10^{-1}$ $2\cdot 10^{3}$ $200f_{\mathrm{dec}}$ $\eta_{\mathrm{link}}^{*}(f_{\eta})$
@@
$p_{\text{g}}$ $0.01$ Threshold
$p_{\text{m}}$ $0.01$
@@ V.1 Current state-of-the-art parameter set
Let us first consider a parameter set inspired by state-of-the-art NV center hardware (see the first column of Table 1). The operation times in this set are based on typical time scales in nitrogen-vacancy centers with a natural 13C concentration Bradley2019 ; pompiliRealizationMultinodeQuantum2021 ; bradleyRobustQuantumnetworkMemory2022 —see App. A for more details. The Bell pair parameter values are a collection of the best parameters in current NV center literature—see the first column of Table 1 for relevant citations. In the following, we explicitly refer to this parameter set as the “state-of-the-art” parameter set. As discussed in more detail in App. A, for this set, the single-click entanglement protocol outperforms the double-click protocol.
@@ V.2.1 Sensitivity to Bell pair fidelity
Firstly, we investigate the influence of the Bell pair fidelity by using a near-future setting parameter set—see the second column in Table 1. Compared to the state-of-the-art parameter set of Sec. V.1, in this set coherence times during entanglement creation and the photon detection probability are one and two orders of magnitude higher, respectively. The double-click entanglement protocol now gives rise to the best combination of entanglement success probability and Bell pair fidelity, as explained in more detail in App. A. This means that these near-future parameters allow for an increase in the link efficiency by four orders of magnitude compared to the state-of-the-art parameter set of Sec. V.1—see Eqs. (1), (3), and (5).
@@
In our Bell pair model, several parameters contribute to the infidelity of the Bell pair states similarly—i.e., through the parameter $\phi$ of Eq. (4) that captures all dephasing noise of the model. To investigate the sensitivity of the performance with respect to the Bell pair fidelity, we vary the influence of dephasing by scaling the probability of double excitation probability and off-resonant excitation errors. These are considered one of the leading error sources in present experiments. hermansQubitTeleportationNonneighbouring2022 We show the results in Fig. 6. In this figure, the bottom of the horizontal axis indicates the Bell pair fidelity; the top indicates the corresponding excitation error probability.